If the initial allele frequencies are p = 0.9 and q = 0.1 and the a allele is a lethal recessive, what will be the frequency of the a allele in the future?
Use the following equation: qn = q0/(1 + nq0), where qn = the new gene frequency, q0 = the initial gene frequency, and n = the number of generations.

1. What will be the frequency of the a allele after 5 generations?

Round your answer to four decimal places.

2.

What will be the frequency of the a allele after 25 generations?

Round your answer to four decimal places.

3. What will be the frequency of the a allele after 100 generations?

Round your answer to four decimal places.
Answer #1

1. After 5 generations according to given formula, qn=q0/(1+nq0) = 0.1/(1+0.5) = 0.0667

2. After 25 generations according to given formula, qn=q0/(1+nq0) = 0.1/(1+2.5) = 0.0286

3. After 100 generations according to given formula, qn=q0/(1+nq0) = 0.1/(1+10) = 0.0091